OpenStudy (anonymous):

show work for -10x-21=-171

OpenStudy (studygurl14):

You need to isolate the x. -10x - 21 = -171 + 21 + 21 __________________ -10x = -171 + 21 What's -171 + 21? And then, do you know what to do after that?

OpenStudy (anonymous):

192/10

OpenStudy (anonymous):

171-21=150/10=15 x=15

OpenStudy (studygurl14):

Not quite. Notice the - in front of the 171. \(\large -171+21\neq 192\)

OpenStudy (anonymous):

-171+21=150/10=15 x=15

OpenStudy (studygurl14):

Oh, right. Nice job